OpenStudy (anonymous):

Trigonometric Integrals \[\int\limits_{}^{}\cos ^{2}\theta \sin \theta \]

hartnn (hartnn):

put u = cos thetha

hartnn (hartnn):

du = ...?

OpenStudy (anonymous):

-sin t

hartnn (hartnn):

du = -sin t dt so, -du = sin t dt

hartnn (hartnn):

integral u^2 (-du) piece of cake :)

OpenStudy (anonymous):

I'm sorry I meant. \[\int\limits_{}^{}\cos ^{2}\theta \sin2\theta \] Is it still the same process Or is it different now

hartnn (hartnn):

write sin 2t = 2 sint cost then put u= sin t

hartnn (hartnn):

sorry, u= cos t

OpenStudy (anonymous):

so iis it -1/2cos^4t

hartnn (hartnn):

-2 integral u^3du -2/4 u^4 +c -1/2 (cos^4 t) +c yup, correct :)

OpenStudy (anonymous):

thanks so much for your help

hartnn (hartnn):

welcome ^_^